# Constrained Denoising, Empirical Bayes, and Optimal Transport

### v1

@@ Appendix E Proofs from Section 4
$\begin{split}&2\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left\langle\nabla\phi^{\ast}(\bar{\eta}),\bar{\eta}-\hat{\eta}\right\rangle\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &\leq 2\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left(\phi^{\ast}(\bar{\eta})-\phi^{\ast}(\hat{\eta})+\frac{1}{2\lambda}\|\bar{\eta}-\hat{\eta}\|^{2}\right)\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &=2\int_{\mathbb{R}^{m}}\phi^{\ast}(\bar{\eta})\textnormal{d}\bar{G}_{n}(\bar{\eta})-2\int_{\mathbb{R}^{m}}\phi^{\ast}(\hat{\eta})\textnormal{d}\hat{G}_{n}(\bar{\eta})+\frac{1}{\lambda}W_{2}^{2}(\hat{G}_{n},\hat{G}_{n})\end{split}$ (E.13)
@@
$\begin{split}&2\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left\langle-\bar{\eta},\bar{\eta}-\hat{\eta}\right\rangle\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &=\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left(\|\hat{\eta}\|^{2}-\|\hat{\eta}-\bar{\eta}\|^{2}-\|\bar{\eta}\|^{2}\right)\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &=\int_{\mathbb{R}^{m}}\|\hat{\eta}\|^{2}\textnormal{d}\hat{G}_{n}(\hat{\eta})-\int_{\mathbb{R}^{m}}\|\bar{\eta}\|^{2}\textnormal{d}\bar{G}_{n}(\bar{\eta})-W_{2}^{2}(\bar{G}_{n},\hat{G}_{n}),\end{split}$ (E.14)
and then plug in (E.13) and (E.14) into (E.12), yielding
$\begin{split}W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\hat{G}_{n}\right)&\leq W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\bar{G}_{n}\right)+\frac{1}{\lambda}W_{2}^{2}(\bar{G}_{n},\hat{G}_{n})\\ &\quad+\int_{\mathbb{R}^{m}}\left(2\phi^{\ast}(\bar{\eta})-\|\bar{\eta}\|^{2}\right)\textnormal{d}\bar{G}_{n}(\bar{\eta})-\int_{\mathbb{R}^{m}}\left(2\phi^{\ast}(\hat{\eta})-\|\hat{\eta}\|^{2}\right)\textnormal{d}\hat{G}_{n}(\bar{\eta}).\end{split}$ (E.15)
@@
$\begin{split}\frac{1}{2L}\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)&\leq\frac{1}{2\lambda}W_{2}^{2}(\hat{G}_{n},\bar{G}_{n})\\ &-\int_{\mathbb{R}^{d}}\left\langle\delta_{\mathcal{B}}(z)-\hat{\delta}_{\mathcal{B}}(z),\delta_{\mathcal{DCB}}(z)\right\rangle\textnormal{d}\bar{F}_{n}(z)\\ &+\frac{1}{2}\int_{\mathbb{R}^{d}}\left\|\delta_{\mathcal{B}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)-\frac{1}{2}\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{B}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z).\end{split}$ (E.16)
It is useful to compare (E.16) with Slawski and Sen [2024, Theorem 7] which derives a similar bound, but with only the first term; the additional terms are due to the fact that $\hat{\delta}_{\mathcal{DCB}}$ comes from an transport problem whose cost function has been estimated.
In the rest of the proof, we will show that each of the terms on the right side of (E.16) is $O_{\mathbb{P}}(\alpha_{n}^{\nicefrac{{1}}{{2}}}\vee\beta_{n}\vee\gamma_{n})$. The first term is the easiest to handle since by the triangle inequality we can bound
@@
E.3 Proofs from Subsection 4.3
Lastly, we give the proof of Theorem 4.7 on the rate of convergence for the general-constrained EB denoiser $\hat{\delta}_{\mathcal{GCB}}$ from Algorithm 3.
@@
In fact, we will prove a slightly stronger claim. For convenience, write $\hat{\pi}_{n}$ for $\hat{\pi}_{\mathcal{GCB}}$, in order to emphasize the dependence on $n\in\mathbb{N}$. Then, we claim that we have $W_{2}(\hat{\pi}_{n},\pi_{\mathcal{GCB}})\to 0$ in probability as $n\to\infty$. Since assumptions (ZC), (BC), (C), and (QG) imply (Theorem 3.12) that $\pi_{\mathcal{GCB}}$ is the unique solution to (3.5), it suffices to show that $\{\hat{\pi}_{n}\}_{n\in\mathbb{N}}$ is $W_{2}$-precompact and that every subsequential $W_{2}$-limit is a solution to (3.5). To see pre-compactness, simply note that the marginal distributions of $\{\hat{\pi}_{n}\}_{n\in\mathbb{N}}$ are $\{\bar{F}_{n}\}_{n\in\mathbb{N}}$ and $\{\hat{H}_{n}\}_{n\in\mathbb{N}}$; the former converges in $W_{2}$ to $F$ hence it is $W_{2}$-pre-compact, and the latter is $W_{2}$-pre-compact since these probability measures have uniformly bounded support.
@@
$\left|\int_{\mathbb{R}^{d}\times K}c_{G}\,\textnormal{d}\pi_{n}-\int_{\mathbb{R}^{d}\times K}\hat{c}_{n}\textnormal{d}\pi_{n}\right|\to 0.$ (E.17)
@@
As before, the first term vanishes because of (EBQ) and the second term is bounded because of (CS). Thus, we have shown (E.17).
Lastly, we show that $\pi_{\infty}$ is in fact optimal for (3.5). To do this, take an arbitrary $\pi\in\Gamma(F;H)$ and use Lemma E.2 to get some $\{\pi_{k}\}_{k\in\mathbb{N}}$ with $\pi_{k}\in\Gamma(\bar{F}_{n_{k}};\hat{H}_{n_{k}})$ and $W_{2}(\pi_{k},\pi)\to 0$ as $k\to\infty$. Then use the Portmanteau lemma, the estimate (E.17), the optimality of $\{\hat{\pi}_{n_{k}}\}_{k\in\mathbb{N}}$, and the estimate (E.17) again, to get:

### v2

@@ Appendix G Proofs from Section 4
$\begin{split}&2\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left\langle\nabla\phi^{\ast}(\bar{\eta}),\bar{\eta}-\hat{\eta}\right\rangle\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &\leq 2\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left(\phi^{\ast}(\bar{\eta})-\phi^{\ast}(\hat{\eta})+\frac{1}{2\lambda}\|\bar{\eta}-\hat{\eta}\|^{2}\right)\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &=2\int_{\mathbb{R}^{m}}\phi^{\ast}(\bar{\eta})\textnormal{d}\bar{G}_{n}(\bar{\eta})-2\int_{\mathbb{R}^{m}}\phi^{\ast}(\hat{\eta})\textnormal{d}\hat{G}_{n}(\bar{\eta})+\frac{1}{\lambda}W_{2}^{2}(\hat{G}_{n},\hat{G}_{n})\end{split}$ (G.13)
@@
$\begin{split}&2\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left\langle-\bar{\eta},\bar{\eta}-\hat{\eta}\right\rangle\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &=\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\left(\|\hat{\eta}\|^{2}-\|\hat{\eta}-\bar{\eta}\|^{2}-\|\bar{\eta}\|^{2}\right)\textnormal{d}\rho_{n}(\bar{\eta},\hat{\eta})\\ &=\int_{\mathbb{R}^{m}}\|\hat{\eta}\|^{2}\textnormal{d}\hat{G}_{n}(\hat{\eta})-\int_{\mathbb{R}^{m}}\|\bar{\eta}\|^{2}\textnormal{d}\bar{G}_{n}(\bar{\eta})-W_{2}^{2}(\bar{G}_{n},\hat{G}_{n}),\end{split}$ (G.14)
and then plug in (G.13) and (G.14) into (G.12), yielding
$\begin{split}W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\hat{G}_{n}\right)&\leq W_{2}^{2}\left((\delta_{\mathcal{B}})_{\#}\bar{F}_{n},\bar{G}_{n}\right)+\frac{1}{\lambda}W_{2}^{2}(\bar{G}_{n},\hat{G}_{n})\\ &\quad+\int_{\mathbb{R}^{m}}\left(2\phi^{\ast}(\bar{\eta})-\|\bar{\eta}\|^{2}\right)\textnormal{d}\bar{G}_{n}(\bar{\eta})-\int_{\mathbb{R}^{m}}\left(2\phi^{\ast}(\hat{\eta})-\|\hat{\eta}\|^{2}\right)\textnormal{d}\hat{G}_{n}(\bar{\eta}).\end{split}$ (G.15)
@@
$\begin{split}\frac{1}{2L}\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{DCB}}(z)-\delta_{\mathcal{DCB}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)&\leq\frac{1}{2\lambda}W_{2}^{2}(\hat{G}_{n},\bar{G}_{n})\\ &-\int_{\mathbb{R}^{d}}\left\langle\delta_{\mathcal{B}}(z)-\hat{\delta}_{\mathcal{B}}(z),\delta_{\mathcal{DCB}}(z)\right\rangle\textnormal{d}\bar{F}_{n}(z)\\ &+\frac{1}{2}\int_{\mathbb{R}^{d}}\left\|\delta_{\mathcal{B}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z)-\frac{1}{2}\int_{\mathbb{R}^{d}}\left\|\hat{\delta}_{\mathcal{B}}(z)\right\|^{2}\textnormal{d}\bar{F}_{n}(z).\end{split}$ (G.16)
It is useful to compare (G.16) with Slawski and Sen [2024, Theorem 7] which derives a similar bound, but with only the first term; the additional terms are due to the fact that $\hat{\delta}_{\mathcal{DCB}}$ comes from an transport problem whose cost function has been estimated.
In the rest of the proof, we will show that each of the terms on the right side of (G.16) is $O_{\mathbb{P}}(\alpha_{n}^{\nicefrac{{1}}{{2}}}\vee\beta_{n}\vee\gamma_{n})$. The first term is the easiest to handle since by the triangle inequality we can bound
@@
G.3 Proofs from Subsection 4.3
Lastly, we give the proof of Theorem C.5 on the rate of convergence for the general-constrained EB denoiser $\hat{\delta}_{\mathcal{GCB}}$ from Algorithm 3.
@@
In fact, we will prove a slightly stronger claim. For convenience, write $\hat{\pi}_{n}$ for $\hat{\pi}_{\mathcal{GCB}}$, in order to emphasize the dependence on $n\in\mathbb{N}$. Then, we claim that we have $W_{2}(\hat{\pi}_{n},\pi_{\mathcal{GCB}})\to 0$ in probability as $n\to\infty$. Since assumptions (ZC), (BC), (C), and (QG) imply (Theorem C.3) that $\pi_{\mathcal{GCB}}$ is the unique solution to (3.4), it suffices to show that $\{\hat{\pi}_{n}\}_{n\in\mathbb{N}}$ is $W_{2}$-precompact and that every subsequential $W_{2}$-limit is a solution to (3.4). To see pre-compactness, simply note that the marginal distributions of $\{\hat{\pi}_{n}\}_{n\in\mathbb{N}}$ are $\{\bar{F}_{n}\}_{n\in\mathbb{N}}$ and $\{\hat{H}_{n}\}_{n\in\mathbb{N}}$; the former converges in $W_{2}$ to $F$ hence it is $W_{2}$-pre-compact, and the latter is $W_{2}$-pre-compact since these probability measures have uniformly bounded support.
@@
$\left|\int_{\mathbb{R}^{d}\times K}c_{G}\,\textnormal{d}\pi_{n}-\int_{\mathbb{R}^{d}\times K}\hat{c}_{n}\textnormal{d}\pi_{n}\right|\to 0.$ (G.17)
@@
As before, the first term vanishes because of (EBQ) and the second term is bounded because of (CS). Thus, we have shown (G.17).
Lastly, we show that $\pi_{\infty}$ is in fact optimal for (3.4). To do this, take an arbitrary $\pi\in\Gamma(F;H)$ and use Lemma G.2 to get some $\{\pi_{k}\}_{k\in\mathbb{N}}$ with $\pi_{k}\in\Gamma(\bar{F}_{n_{k}};\hat{H}_{n_{k}})$ and $W_{2}(\pi_{k},\pi)\to 0$ as $k\to\infty$. Then use the Portmanteau lemma, the estimate (G.17), the optimality of $\{\hat{\pi}_{n_{k}}\}_{k\in\mathbb{N}}$, and the estimate (G.17) again, to get: